# Unveiling Transformer Perception by Exploring Input Manifolds

### v1

@@ 3.2 Interpretability
Algorithms 1 and 2 allow for the exploration of the equivalence classes in the input space of a Transformer model. However, the points explored by these algorithms may not be directly interpretable by a human perspective. For instance, an image or a piece of text may need to be decoded to be “readable” by a human observer. Furthermore, we present an interpretation of the eigenvalues of the pullback metric which allows us to define a feature importance metric. We present two interpretability methods for Transformers based on input space exploration. Both methods are then demonstrated on a Vision Transformer (ViT) trained for digit classification [8], and two BERT models, one trained for hate speech classification and the other trained for MLM [7, 19].
Algorithm 3 Feature Importance Analysis Using Pullback Metric $g^{0}_{x^{e}}$
1:Inputs:
2:      Transformer model $T$ with: Tokenizer $t_{T}$, Embedding layer $e_{T}$, Intermediate layers $l_{T}$
3:      Input data $x$
4:Tokenize input $x$ to obtain tokens $x^{t}=t_{T}(x)$
5:Compute embeddings $x^{e}=e_{T}(x^{t})$
6:Compute intermediate representations $g_{l_{T}(x^{e})}^{n}$
7:Calculate the pullback metric $g^{0}_{x^{e}}$
8:Diagonalize $g^{0}_{x^{e}}$ to extract eigenvalues
9:Identify the maximum eigenvalue for each embedding, indicating its importance
10:Output: Heatmap of embedding importance based on the eigenvalues
Algorithm 4 Exploration of Embedding Space in Transformers
1:Inputs:
2:      Transformer model $T$ with: Tokenizer $t_{T}$, Embedding layer $e_{T}$, Intermediate layers $l_{T}$
3:      Input data $x$ (image or text)
4:Retrieve segments $x^{t}=t_{T}(x)$.
5:Choose segments $P=\{p|p\in x^{t}\}$ for updates; keep others unchanged.
6:Compute embeddings $x^{e}=e_{T}(x^{t})$.
7:Apply SiMEC or SiMExp on $x^{e}$, updating embeddings for segments in $P$.
8:Outputs: Modified input embedding, one for each SiMEC/SiMExp iteration.
Feature importance.
Consider a Transformer model $T$ whose architecture includes a tokenizer $t_{T}$ (or patcher for images) that segments the input so that each segment can be converted into a continuous representation by an embedding layer $e_{T}$. This results in a matrix of dimensions $n_{s}\times h$, where $n_{s}$ represents the number of segments, and $h$ denotes the hidden size of the model’s embeddings. The eigenvalues of the pullback metric can be used to deduce the importance of each embedding and, by extension, the significance of the segments they represent, with respect to the final prediction. The process for determining the importance of textual tokens or image patches is outlined in Algorithm 3. The appearance of the resulting heatmaps varies according to the type of input used. An example of experiments with ViT on the MNIST dataset [12] is shown in Figure 1 that depicts heatmaps for two MNIST instances.
Figure 1: Example output from Algorithm 3 applied to digit classification. These two instances are predicted as 3 (left) and 4 (right). The brightness of the color indicates the eigenvalue’s magnitude. The brighter the color, the more sensitive the patch. This indicates that changes in the values of these sensitive patches are likely to have a greater impact on the prediction probabilities. Each patch in the heatmap corresponds to a $2\times 2$ square pixel.
Figure 2, on the left, illustrates two experiment using Algorithm 3 on both a BERT model for hate speech detection and a BERT model for MLM.
Figure 2: Example outputs from Algorithm 3. The darker the color, the higher the token’s eigenvalue. Left: The sentence analysed is classified as “offensive” by the BERT for hate speech detection, with significant contributions from tokens [CLS], politicians, corrupt, and ##eit (part of the word deceitful). Right: Example instance processed by a BERT model for masked language modeling. [MASK] is predicted as “ham”, with the most influential tokens being pizza and cheese.
Interpretation of input space exploration.
@@
Using SiMEC and SiMExp to explore the embedding space reveals how Transformer models perceive equivalence among different data points. Specifically, these methodologies facilitate the sequential acquisition of embedding matrices $p_{0}\dots p_{K}$ at each iteration, as detailed in Algorithms 1 and 2. Algorithm 4 implements a practical application of the SiMEC/SiMExp approach with Transformer models. A key feature of this method is its ability to selectively update specific tokens (for text inputs) or patches (for image inputs) during each iteration. This selective updating allows us to explore targeted modifications that prompt the model to either categorize different inputs as the same class or recognize them as distinct. Unlike traditional approaches where modifications are predetermined, this method lets the model itself guide us to understand which data points belong to specific equivalence classes. To interpret embeddings resulted from the exploration process, they must be mapped back into a human-understandable form, such as text or images. The interpretation of an embedding vector depends on the operations performed by the Transformer’s embedding module $e_{T}$. If $e_{T}$ consists only of invertible operations, it is feasible to construct a layer that performs the inverse operation relative to $e_{T}$. The output can then be visualized and directly interpreted by humans, allowing for a comparison with the original input to discern how differences in embeddings reflect differences in their representations (e.g., text, images). If the operations in $e_{T}$ are non-invertible, a trained decoder is required to reconstruct an interpretable output from each embedding matrix $p_{0}\dots p_{K}$. When using a BERT model, it is feasible to utilize layers that are specialized for the masked language modeling (MLM) task to map input embeddings back to tokens. This approach is effective whether the BERT model in question is specifically designed for MLM or for sentence classification. In the case of sentence classification models, it is necessary to select a corresponding MLM BERT model that shares the same internal architecture, including the number of layers and embedding size.
@@
Algorithm 5 depicts the process of interpreting Algorithm 4 outputs for both ViT and BERT experiments. After initializing the decoder according to the model type, the embeddings $p_{0}\dots p_{K}$ need to be constrained to a feasible region. This region is defined by the distribution of embeddings derived from the original input instances. Next, the embeddings are decoded, and the selected segments for exploration are extracted. These segments are then used to replace the corresponding parts of the original input instance.
@@
2:      Transformer model $T$ with: Tokenizer $t_{T}$, Embedding layer $e_{T}$, Intermediate layers $l_{T}$
3:      Modified embeddings $p_{0}\dots p_{K}$ resulted from Algorithm 4 applied on an input $x$
4:      $P=\{p|p\in x^{t}\}$ indices of updated segments
@@
6:      Initialize decoder $d$ with weights from $e_{T}$.
@@
9:Compute embeddings distributions for original input data
10:Use the original embeddings distributions to cap $p_{0}\dots p_{K}$
11:Decode modified embeddings $p_{0}\dots p_{K}$ using $d$ to generate the corresponding images/sentences $X^{\prime}=x^{\prime}_{0}\dots x^{\prime}_{K}$.
12:For each $x^{\prime}\in X^{\prime}$: replace segments relative to indices $P$ in $x$ with those in $x^{\prime}$.
13:Outputs:
14:Modified input images/sentences, one for each SiMEC/SiMExp iteration.
Figure 3 depicts an example outcome of Algorithm 5 applied on a ViT exploration experiment.
Figure 3: Example of SiMEC and SiMExp output interpretation for ViT digit classification. Left: Original MNIST image of an “8”. Center: Interpretation of a $p_{1000}$ from a SiMEC experiment, where $p_{1000}$ is predicted as “8”. Right: Interpretation of a $p_{1000}$ from a SiMExp experiment, where $p_{1000}$ is predicted as “4”. All patches are subject to SiMEC and SiMExp updates.
Given that the interpretation process includes both a capping step and a decoding step (lines 10 and 11 of Algorithm 5), it’s important to note that there isn’t a direct 1:1 correspondence between each iteration’s update and the interpretation outcomes. Our primary focus is on exploring the input embedding space, rather than the input image or input sentence spaces. For further investigation, we provide a detailed discussion on considering interpretation outputs as alternative prompts in Section 4.
@@ 4 Experiments
Experiments are conducted on textual and visual data. We aim to perform a preliminary investigation of 3 features of our approach: (i) how the class probability changes on the decoded output of SiMEC/SiMExp, (ii) what is the trade-off between the quantity and the quality of the output, and (iii) how our method can be used to extract feature importance-based explanations.
In the textual case, we experiment with hate speech classification datasets: we use HateXplain [^2] MIT License [13], which provides a ground truth for feature importance, plus a sample of 100 hate speech sentences generated by prompting ChatGPT [^3] Used prompts are included in the Supplementary Materials., which serve purposes (i) and (ii). In the visual case, we perform experiments on MNIST [12] dataset.
Using interpretation outputs as alternative prompts
An interesting investigation is to determine if our interpretation algorithm (Algorithm 5) can generate alternative prompts that stay in the same equivalence class as the original input data or move to a different one, based on SiMEC and SiMExp explorations. We test how the probability assigned to the original equivalence class by the Transformer model changes as the SiMEC and SiMExp algorithms explore the input embedding manifold.
For BERT experiments we generate prompts to inspect the probability distribution over the vocabulary for tokens updated by Algorithms 1 and 2. We decode the updated $p_{0}\dots p_{K}$ using Algorithm 5, focusing on tokens updated through the iterations. For each of these decoded tokens, we extract the top-5 scores to obtain 5 alternative tokens to replace the original ones, creating 5 alternate prompts. We then extract the prediction $i^{*}=\operatorname*{arg\,max}_{i}{y_{i}}$ for the original sentence, which represents the output whose equivalence class we aim to explore. Finally, we classify the new prompts, obtaining the corresponding predictions $Y=\mathbf{y}^{(0)}\dots\mathbf{y}^{(K)}$, where each $\mathbf{y}^{(k)}\in\mathbb{R}^{N}$, $N$ being the number of prediction classes. We visualize the prediction trend for the $i^{*}$th value in every $\mathbf{y}^{(0)}\dots\mathbf{y}^{(K)}$ categorizing the images into two subsets: those that lead to a change in prediction $Y_{c}=\{\mathbf{y}^{(k)}\in Y\mid\operatorname*{arg\,max}_{i}{y^{(k)}_{i}}\neq i^{*}\}$ and those that don’t $Y_{s}=\{\mathbf{y}_{i}\in Y\mid\operatorname*{arg\,max}_{i}{y^{(k)}_{i}}=i^{*}\}$.
Figure 4: Analysis involving results SiMEC and SiMExp applied to BERT for hate speech detection. Left: Prediction values for $i^{*}$ for each $\mathbf{y}\in Y_{c}$. Right: Prediction values for $\mathbf{y}\in Y_{s}$.
Sentence classification experiments [^4] Model used: huggingface.co/ctoraman/hate-speech-bert involved 1000 iterations from both SiMEC and SiMExp, applied to a subset of 8 sentences from the ChatGPT hate speech dataset. The plot on the left side of Figure 4 illustrates that, as the original embeddings are increasingly modified, SiMExp tends to produce alternatives with lower prediction values for $i^{*}$ compared to SiMEC. Thus, even if predictions change in SiMEC experiments, the equivalence class prediction value remains approximately constant and higher than in SiMExp. Considering the plot on the right side of Figure 4, SiMExp identifies prompts that lower the prediction value for $i^{*}$. ViT and MLM experiments are detailed in the Supplementary Materials.
@@
We measure the time required to explore the input space of a ViT with the SiMEC algorithm and compare it with a perturbation-based method. The perturbation-based method mimics a trial-and-error approach as it takes an input image and, at each iteration, perturbs it by a semi-random vector $\mathbf{v}_{t+1}=a_{t}\mathbf{v}_{t}+\eta\epsilon$, where $a_{t}=1$ if $y_{t}=y_{t-1}$, $a_{t}=-1$ otherwise, $\epsilon$ is an orthogonal random vector from a standard normal distribution and $\eta$ is the step length. With the perturbation, we obtain a new image, then check whether the model yields the same label for the new image. The perturbation vector is re-initialized at random from a normal distribution $20\%$ of the times to allow for exploration. We construct this method to have a direct comparison with ours in the absence of a consolidated literature about the task.
We train a ViT model having 4 layers and 4 heads per layer on the MNIST dataset [^5] Using Adam optimizer, the model achieved the highest validation accuracy ($96.25\%$) in 20 epochs.. The SiMEC algorithm is run for 1000 iterations, so that it can generate 1000 examples starting from a single image. In a sample of 100 images, the average time is approximately 339 seconds. [^6] All experiments are based on the current PyTorch implementation of the algorithms and run on a Ubuntu 20.04 machine endowed with one NVIDIA A100 GPU and CUDA 12.4. In the same time, the perturbation-based algorithm can produce up to 36000 images. However, we notice that the perturbation-based algorithm ends up producing monochrome (pixel color has zero variance) or totally noisy images, which provide little information about the behavior of the model. Excluding only the images with low color variance ($<0.01$), we are left, on average, with $19$ images (standard deviation $13.9$). SiMEC, in contrast, doesn’t present this behavior, as all 1000 images have high enough intensity variance and are thus useful for explainability purposes.
As BERT has many more parameters with respect to our ViT model, processing textual data takes longer. Specifically, in a sample of 16 sentences, the average time needed to run 1000 iterations on a sentence is 7089 seconds, taking into account both MLM and classification experiments.
Feature importance-based explanations
We compare our method against Attention Rollout (AR) [1] and the Relevancy method proposed by  Chefer et al. [6]. In the textual case, we provide a quantitative evaluation using the HateXplain dataset, which contains 20147 sentences (of which 1924 in the test set) annotated with normal, offensive and hate speech labels as well as the positions of words that support the label decision. We then measure the cosine similarity between the importance assigned by each method to each word in a sentence and the ground truth. Notice that, since the dataset contains multiple annotations, the ground truth $y$ for each word $w$ is obtained as the average of the binary labels assigned by each annotator, and therefore $y(w)\in[0;1]$. We also normalize all scores in $[0;1]$ so to have them on the same scale. The average similarity achieved by our method is 0.707 (standard deviation $\sigma$ = 0.302), against 0.7 ($\sigma$ = 0.315) for Relevancy and 0.583 ($\sigma$ = 0.318) for AR. This proves our method to be more effective in finding the most sensitive tokens for classification. We provide an example on image classification in the Supplementary Materials.

### v2

@@ 3.2 Interpretation
Algorithms 1 and 2 allow for the exploration of the equivalence classes in the input space of a Transformer model. However, the points explored by these algorithms may not be directly interpretable by a human perspective. For instance, an image or a piece of text may need to be decoded to be “readable” by a human observer. Here we present an interpretation method for Transformers based on input space exploration, which is then demonstrated on two Vision Transformer (ViT) models trained for image classification [8], and two BERT models, one trained for masked language modeling (MLM) [7] and the other fine-tuned for text classification [18].
@@
Using SiMEC and SiMExp to explore the embedding space reveals how Transformer models perceive equivalence among different data points. Specifically, these methodologies facilitate the sequential acquisition of embedding matrices $x^{(0)}\dots x^{(K)}$, as detailed in Algorithms 1 and 2. A key feature of the SiMEC/SiMExp approach is its ability to selectively update specific tokens (for text inputs) or patches (for image inputs) during each iteration. This selective update allows to explore targeted modifications that prompt the model to either categorize different inputs as the same class or recognize them as distinct. Unlike other approaches [24, 12] where perturbations are predetermined, this method lets the model itself guide us to understand which data points belong to specific equivalence classes.
To interpret embeddings produced by the exploration process, they must be mapped back into a human-understandable form, such as text or images. The interpretation of an embedding vector depends on the operations performed by the Transformer’s embedding module $\mathcal{E}_{T}$. If $\mathcal{E}_{T}$ consists only of invertible operations, it is feasible to construct a layer that performs the inverse operation relative to $\mathcal{E}_{T}$. The output can then be visualized and directly interpreted by humans, allowing for a comparison with the original input to discern how differences in embeddings reflect differences in their representations (e.g., text, images). If the operations in $\mathcal{E}_{T}$ are non-invertible, a trained decoder is required to reconstruct an interpretable output from each embedding matrix $x^{(1)},\cdots,x^{(K)}$. Such operation injects some unforeseen noise into the interpretation results, which is investigated in the experimental setting. In practice, we construct the ViT models such that the embedding layer is invertible, whereas for BERT models it is feasible to exploit layers that are specialized for the MLM task to map input embeddings back to tokens. This approach is effective whether the BERT model in question is specifically designed for MLM or for sentence classification. In the case of sentence classification models, it is necessary to select a corresponding MLM BERT model that shares the same internal architecture, including the number of layers and embedding size.
@@
Algorithm 3 depicts the process of interpreting SiMEC/SiMExp outputs for both ViT and BERT experiments. After initializing the decoder according to the model type, the embeddings $x^{(1)},\cdots,x^{(K)}$ are decoded and the selected segments for exploration are extracted. These segments are then used to replace the corresponding parts of the original input instance.
@@
2:   Transformer model $T$ with: Patcher/Tokenizer $\mathcal{T}_{T}$, Embedding layer $\mathcal{E}_{T}$. Input image/text $z$
3:   Modified embeddings $x^{(1)}\dots x^{(K)}$ resulted from Algorithm 1 or  2 applied on $x^{(0)}=\mathcal{E}_{T}(\mathcal{T}_{T}(z))$
4:   $P\subseteq\{1,\dots,dim(z)\}$ indices of patches/tokens to update
@@
6:   Initialize decoder $d$ with weights from $\mathcal{E}_{T}$.
@@
9:Decode modified embeddings $x^{(0)}\dots x^{(K)}$ using $d$ to generate the images/sentences $Z^{\prime}=Z^{\prime}_{0}\dots Z^{\prime}_{K}$.
10:For each $z^{\prime}\in Z^{\prime}$: replace segments relative to indices $P$ in $z$ with those in $z^{\prime}$.
11:Outputs:
12:Modified input images/sentences, one for each SiMEC/SiMExp iteration.
Figure 1 (top left) presents the outcome of applying Algorithm 3 to a ViT exploration experiment on a CIFAR10 image. Both SiMEC and SiMExp produce visually similar outputs—each still resembling a “cat” to a human observer—yet the SiMExp interpretation is classified as “dog” at iteration 750. This demonstrates how subtle modifications, such as changes in background pixels, can significantly influence model predictions, even when such changes are perceived as irrelevant by humans, as also noted in [24]. A clear difference emerges in the exploration dynamics of the two algorithms (Figure 1, top right): SiMExp progresses in a more straight and directed manner, reflecting its goal of escaping the initial equivalence class. This divergence is further illustrated in the bottom right subplots, where the class probability distributions remain stable during SiMEC exploration but show notable fluctuations under SiMExp exploration. The lower part of Figure 1 shows a similar example for textual data from the Measuring-Hate-Speech (MHS) dataset. In this case, SiMExp identifies an alternative sentence that is classified as “Hatespeech”, contrasting with the original input, which had been classified as “Offensive”.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 1: (Top figure) Example of exploration on a CIFAR10 image using SiMEC and SiMExp. Left: Original image, followed by interpretation outputs of $x_{750}$ from SiMEC (middle) and SiMExp (bottom). Right top: SVD projection of the explored points $x^{(1)},\cdots,x^{(K)}$ for SiMEC (red) and SiMExp (blue), where color intensity encodes iteration progress (darker colors correspond to later iterations), and point shapes indicate predicted class labels. Right bottom: Evolution of class probabilities over iterations, for SiMEC (left) and SiMExp (right). (Bottom figure) Example of exploration on an MHS sentence using SiMEC and SiMExp. Visualization layout and interpretation are analogous to the top figure.
@@ 4 Experiments
Experiments are conducted on textual and visual data and are aimed at two objectives: (i) obtaining an empirical verification of the behavior of SiMEC and SiMExp under diverse settings, and (ii) verifying the consistency of interpretation outputs with the ones from exploration only, in order to test their usability as alternative input data.
In each data modality, we experiment with two datasets presenting different features: (i) MNIST [13], a grayscale digit image dataset; (ii) CIFAR10 [11], a RGB object image dataset; (iii) WinoBias [25], a textual dataset for MLM, especially focused on gender bias; (iv) Measuring-Hate-Speech (MHS) [16], a textual dataset for Text Classification, especially focused on hate speech detection. We trained one ViT model for each image dataset, and we used pretrained BERT models for MHS and WinoBias. More details about adopted models, experimental results in further configurations, and full experimental details are provided in the Supplementary Materials.
@@
For objective (i) we consider the following metrics: effectiveness of the exploration can be measured by changes in prediction probabilities as well as estimation of the hyper-volume explored, while speed is assessed in terms of total time (in seconds) [^2] All experiments are based on the current PyTorch implementation of the algorithms and run on a Ubuntu 22.04 machine endowed with one NVIDIA H100 GPU and CUDA 12.4.. Algorithms are run for $K=1000$ iterations, which we prove sufficient to capture their behavior, with delta multiplier $\eta\in\{1;10\}$, the latter used with the aim to verify whether it is possible to speed up the process pace without compromising its stability. Finally, the experiments reported here all refer to the configuration in which all patches of an image are modified, while for textual inputs only the token with the highest attribution value is subject to exploration.
Given the predictions obtained by re-applying the models’ encoder and classifier layers to the modified embeddings $x^{k}$ at each iteration $k$, we observe the changes in class probabilities. The theoretical results suggest that SiMEC should induce minimal fluctuations in them, while SiMExp should yield rapidly changing probabilities, up to prediction changes.
(a) SiMEC and SiMExp
(b) Baseline gradient-based Gaussian noise approaches
Figure 2: Mean and standard deviation (where applicable) of probability values for the original class (solid line) and the top predicted class (dashed line) based on embeddings obtained during exploration, across iterations and datasets. Subfigure (a) depicts the behavior of SiMEC (orange) and SiMExp (blue), while subfigure (b) reports the behavior of corresponding baseline algorithms. SiMExp results in a notable decrease in the probability of the original class, while the probability of the highest-scoring class decreases to a lesser extent, indicating a shift in the most probable class.
Figure 2(a) depicts the empirical reflection of the theoretical results. Focusing on the probability of the original input class (i.e. class predicted at $k=0$), we see that SiMEC manages to keep it constant while SiMExp makes it drop significantly within the first 1000 iterations. As a baseline, we compare our results with a gradient-based Gaussian-noise approach which updates the input embedding $x^{(k)}$ at each iteration by $\pm\overline{\delta}\nabla\mathcal{N}(x^{(k)})+\epsilon$, where the sign is determined by what exploration we are performing (same class vs other class) and $\epsilon$ is a small Gaussian noise vector orthogonal to the gradient so to guarantee exploration. Applying the same number of iterations and the same average step size $\overline{\delta}$ in each experiment allows us to conclude that SiMEC and SiMExp are significantly more effective than the baselines for staying in the original equivalence class and moving to another class respectively. Indeed, in SiMEC original class probability always follows strictly the top class probability, which in the baseline is rarely the case; in SiMExp the two probabilities tend to diverge as predicted by the theoretical analysis, while in the baseline they remain close to one another.
In order to verify the actual shift in class probabilities, we report in Table 1 the statistics about average class ranking changes, which indicate a clear tendency of SiMExp changing equivalence classes compared to SiMEC.
MNIST CIFAR10 WinoBias MHS
SiMEC $\eta=1$ 0.03(0.18)0.03\,(0.18) $0.0$ $0.0$ $0.0$
$\eta=10$ 0.07(0.25)0.07\,(0.25) 0.12(0.33)0.12\,(0.33) $0.0$ $0.0$
SiMExp $\eta=1$ 1.31(1.64)1.31\,(1.64) 2.60(3.35)2.60\,(3.35)** 0.13(0.33)0.13\,(0.33) 0.75(0.89)0.75\,(0.89)**
$\eta=10$ 11.91(11.87)11.91\,(11.87)** 18.70(15.90)18.70\,(15.90)** 3.25(2.41)3.25\,(2.41)** 3.64(3.43)3.64\,(3.43)**
Table 1: Average class ranking changes per $1000$ iterations across datasets and $\eta$ values. The class ranking change is computed by counting pairwise inversions in class rankings before and after each exploration update. Results are reported as mean (standard deviation) over multiple runs and inputs. The symbol ** indicates that these ranking changes involve at least once a prediction change, i.e., a change in the top predicted class. SiMExp induces substantially more ranking and prediction changes than SiMEC, especially for larger $\eta$.
Furthermore, we estimate the per-patch explored hypervolume by reducing embedding vectors to the first $n$ principal components, retaining 90% of the total variance, and computing at each iteration $k$ the element-wise difference $\Delta^{(k)}_{i}=(\max_{t=0,\dots,k}x^{(t)}_{i}-\min_{t=0,\dots,k}x^{(t)}_{i})^{1/n}$ (the power $1/n$ allows for more stable computation). The product of the components of $\Delta^{(k)}$ gives the volume of an hyper-dimensional cuboid which contains the explored region and is thus an over-estimate of the scope achieved by the exploration. By computing the average volume ratio $\rho_{V}=({\Pi\Delta^{(K)}_{SiMExp}}/{\Pi\Delta^{(K)}_{SiMEC}})^{n}$, we empirically verified that SiMExp explores a portion of space that is bigger than the one explored by SiMEC by an order of $10^{1}$. We validated these results by performing Welch t-tests on $\rho_{V}$: all p-values resulted lower than $10^{-3}$. Thus we conjecture that the exploration took a privileged direction on SiMExp experiments, thus making the volumes increasing faster than in SiMEC experiments.
Finally, we measure the time required to explore the input space of a model with the SiMEC and SiMExp algorithms. Means (and standard deviations) of required times are computed per patch/token and per iteration: $0.126$ s ($0.008$) for CIFAR10, $0.050$ s ($0.004$) for MNIST, $0.300$ s ($0.020$) for Winobias, and $0.310$ s (0.074) for MHS.
Using interpretation outputs as alternative input data
Objective (ii) is to assess whether our interpretation algorithm (Algorithm 3) can generate alternative input data that either preserve the original input’s equivalence class or shift to a different one, depending on the exploration dynamics of SiMEC and SiMExp. The mean difference of pixel/tokens generated from iteration to iteration amounts at $2.219\cdot 10^{-3}$ for SiMEC experiments, and at $83.028\cdot 10^{-3}$ for SiMExp experiments; these values increase as the number of explored patches. These results show that our algorithms generate diverse outputs across iterations, especially when the exploration is performed following the SiMExp algorithm.
Beyond output diversity, we evaluate whether and how the model’s prediction for the original equivalence class evolves as SiMEC and SiMExp explore the embedding space. Specifically, we track how the class probabilities from the modified embeddings change when decoded back into the data domain every $\hat{k}$ iterations. This evaluation differs from simply observing changes in embedding predictions. The projection step (Algorithms 1 and 2, step 11) constrains modifications to an $L^{\infty}$ sphere containing the data domain, not the exact input space. Additionally, the decoder itself introduces approximations, either due to model limitations or numerical errors. These factors can cause discrepancies between predictions from embeddings and from decoded interpretations. To quantify this effect, we compute the average Wasserstein distance ($p=1$) between probability distributions predicted from embeddings and their corresponding interpretation outputs. Wasserstein distance is preferred over KL divergence for its interpretability in this context. Results show that, with a median Wasserstein distance of $0.0$, SiMEC produces interpretation outputs whose predicted probabilities remain very close to those of the embeddings, indicating consistent generation of reliable alternative input data. In contrast, SiMExp’s outputs exhibit larger and more variable Wasserstein distances, whose distribution has a median of $0.049$, highlighting inconsistencies between embeddings and their interpretations.
In cases where predictions on SiMExp’s interpretations initially differ from prediction on SiMExp’s embeddings, an average of $10.9\%$ of these misaligned predictions eventually realign as exploration progresses. On average, this “catch-up” effect occurs after an average of $290.65$ iterations, suggesting that longer exploration trajectories (beyond $1000$ iterations) could further improve alignment. Supporting this observation, we find an average positive Pearson correlation of $0.32$ (average p-value $0.08$) between top class’ probability predicted on SiMExp’s embeddings and its corresponding probability prediction on SiMExp’s interpretations, indicating a trend towards convergence.